A binomial random variable has a mean equal to 181 and a standard deviation of 6.
Find the value of n.
Find the value of p

Answer by stanbon(75887)

A binomial random variable has a mean equal to 181
and a standard deviation of 6.
Find the value of n.
Find the value of p
----
mean = np = 181
std = sqrt(npq) = 6
-----
npq = 6^2 = 36
---
npq/np = 36/181
q = 36/181
Then p = 1-(36/181) = 145/181
---------------
Solve for "n":
np = 181
n = 181/(145/181) = 181^2/145 = 226 when rounded up
-----------
Cheers,
Stan H.
============